OpenStudy (anonymous):

how we know the cube root of 512 is 8

OpenStudy (anonymous):

just see 8*8*8=512 or not if it is then cube root of 512 is 8

OpenStudy (anonymous):

how we know it is 8

OpenStudy (anonymous):

just factorise 512 with the help of 2's after factoring you will find 512=2^9 while 8=2^3 so 512=8^3

OpenStudy (saifoo.khan):

Prime factorization.

OpenStudy (anonymous):

devide 512 by it's prime factors (ie. numbers which can devide 512 and not yield fractions and which themselves have only 2 factors. these are 1,2,3, 5, 7, 11,13,17, etc you can pick out which ones work. and always use the lowest prime factor(LPF) 512=(2*256) lpf=2 256=(2*128) lpf=2 128=(2*64) 64=(2*32) 32=(2*16) 16=(2*8) 8=(2*4) 4=(2*2) 2=(2*1) here the lowest prime factors are 2 everywher in the ladder (Recall that \[8=2^{3}\] so 512=2x2x2x2x2x2x2x2x2= (2x2x2)x(2x2x2)x(2x2x2)=8x8x8 . DONE! Use the ladder method and find the LPF for each step